OpenStudy (mazehaq):

3(x-3)^1/3 - 4(x-3)^4/3 + (x-3)^7/3 = 0 To make it easier, I substituted (x-3) with x, in which I got x^1/3(3 - 4x +x^2) ...but it can't be factored. Help me out! It's a fairly easy question I bet, but I need a nice way of solving it. Thank you! :)

OpenStudy (lgbasallote):

actually you might want to substitute x - 3 with a instead to avoid confusion

OpenStudy (lgbasallote):

but you're right anyway... except for the last part

OpenStudy (lgbasallote):

wait on second thought you're right

OpenStudy (lgbasallote):

and it can be factored... \[\huge a^{\frac 13} (3 - 4a + a^2) \implies a^{\frac 13} (a^2 - 4a + 3)\] can you see it now?

OpenStudy (lgbasallote):

a^2 - 4a + 3 is factorable

OpenStudy (mazehaq):

Oh okay! I kept thinking that it was a^2 + 4a - 3

OpenStudy (lgbasallote):

oh. no wonder

OpenStudy (lgbasallote):

yeah. that one isn't factorable

OpenStudy (mazehaq):

so 3 - 4a + a^2 is a^2 - 4a + 3 ...right?

OpenStudy (lgbasallote):

yes

OpenStudy (mazehaq):

Okie dokie thank you!

OpenStudy (lgbasallote):

welcome

OpenStudy (mazehaq):

So wait, how would I figure out the final answer, I am at (x-3)^1/3 (x-7)(x-7)=0

OpenStudy (lgbasallote):

they shouldn't be equal...

OpenStudy (lgbasallote):

how did you factor a^2 - 4a + 3?

OpenStudy (mazehaq):

I skipped a few steps. Remember a = (x-3) a^1/3 (a-1)(a-3) = 0 (x-3)^1/3 [(x-3-1)(x-3-3)] = 0 (x-3)^1/3 (x-4)(x-6) = 0

OpenStudy (mazehaq):

sorry about my previous comment tho, I totally typed it wrong, my computer is lagging so bad.

OpenStudy (lgbasallote):

right. that's better

OpenStudy (mazehaq):

So how do I solve the end?

OpenStudy (mazehaq):

(x-3)^1/3 (x-4)(x-6) = 0

OpenStudy (lgbasallote):

that s the end

OpenStudy (lgbasallote):

isn't it?

OpenStudy (lgbasallote):

or do you want x?

OpenStudy (mazehaq):

don't I make each = to zero

OpenStudy (mazehaq):

Like, 2 of my answers will be 4 & 6...

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now..just equate (x-3)^1/3 to 0

OpenStudy (mazehaq):

I know but how :(

OpenStudy (lgbasallote):

\[(x - 3)^{\frac 13} = 0\]

OpenStudy (lgbasallote):

take the cube root of both sides..what happens?

OpenStudy (mazehaq):

x-3=0?

OpenStudy (lgbasallote):

yes

OpenStudy (mazehaq):

therefore x = 3?

OpenStudy (lgbasallote):

now solve for x

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so there's your answer

OpenStudy (mazehaq):

Okay, so answer is 3,4,6 :) thanks!

OpenStudy (lgbasallote):

yup and welcome